OpenStudy (anonymous):

I need with a question and can you try to explain it to me because it is confusing me Part A: If (7^2)x = 1, what is the value of x? Explain your answer. (5 points)

OpenStudy (texaschic101):

(7^2)x = 1 49x = 1 x = 1/49 check... (7^2)x = 1 (7^2)(1/49) = 1 (49/1)(1/49) = 49/49 which equals 1

OpenStudy (anonymous):

thank you